Session: June 2008 (session 2)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Hartebeesthoek (Jonathan Quick at Tue Jun 17 14:06:19 2008) Success :
No known problems

Medicina (Roda at Tue Jun 17 15:43:21 2008) Success :

Onsala (Michael Lindqvist at Tue Jun 17 16:37:31 2008) Success :
Source to low at last scans. Mark5 error, 2008.169.14:32:04.89?ERROR m5 -900 Previous recording still active. Oterwise no known problems.

Torun (Kaz Borkowski at Wed Jun 18 07:00:34 2008) Minor failures :
11:08 to 11:17 UT observations unexpectedly switched (probably by a short power loss) to C-band receivers. Also, antenna stopped before the last scan to avoid entering below the telescope horizon.

Effelsberg (Uwe Bach at Wed Jun 18 13:33:17 2008) Minor failures : LATE
Couldn't start before 8 UT because of a hard disk failure on one of the antenna control PCs.

Westerbork (Geert Kuper at Wed Jun 18 14:41:05 2008) Minor failures : RECE
Single dish was used. Please change the coordinates in the vexfile accordingly. At UT 10:00 our TMS system was restarted. It took over the single dish receiver steering with the wrong frequency. It was noticed at 12:29 UT, so we have no data form 10:00-12:29 UT. To prevent this in future, the warning system has been adapted.

Jodrell_Bank (Alastair Gunn at Thu Jun 19 10:49:14 2008) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Thu Jun 19 10:49:37 2008) Success :
No known problems.

Noto (P. Cassaro at Wed Jun 25 10:52:20 2008) Success :